# New Method for Robust Analysis of Critical Behavior of Magnetic Systems

### v1

@@ Abstract
Here, we present new methods for critical analysis to determine the range of exchange interaction(s) and appropriate values of critical exponents for different magnetic systems. From computational and experimental investigations of magnetic behavior of Ni and Gd, we show that the critical behavior remains the same on either side of transition temperature. Using our proposed method of analysis for Gd, we find a critical role of competing interactions where the local electron moments follow 3D Ising type short-range interactions whereas the itinerant electron moments constitute mean-field type long-range RKKY interaction.
@@
In the studies of phase transitions for cases, such as condensation of gases, melting of solids, magnetic ordering in solids, etc., the understanding of the behavior of the systems adjacent to the critical point is of fundamental interest [1, 2, 3]. Generally, the physics of phase transition is governed by the behavior of the microscopic constituents which interact strongly in a cooperative fashion. Critical exponents (CEs), which define the power law behavior of physical quantities exhibiting singularities at the critical points, indicate the qualitative nature, i.e., the critical behavior of a system [4, 5]. The types of cooperative interactions (short-range, long-range, etc.) among the constituent entities play a significant role in defining the CEs. In a magnetic system, the magnetic moments are known to interact typically via different exchange interactions resulting in a set of CEs [6, 7, 8, 9, 10, 11, 12, 13]. Different methodologies have been suggested for determination of a set of CEs ($\alpha,\beta,\gamma$, $\delta$, etc.) corresponding to physical quantities such as specific heat ($C_{\rm{v}}$), spontaneous magnetization ($M$), susceptibility ($\chi$) and magnetization isotherm recorded at the paramagnetic transition temperature ($T_{\rm{C}}$), respectively for various universality classes (UCs), viz., Heisenberg, XY, Ising, mean field (MF), or tricritical MF [4, 5]. The CEs are mutually related and they depend critically on the dimensionality (of space and spin) and the range of microscopic interactions. Although, most methodologies suggest same CEs for two sides of a transition, a recent theoretical work suggested otherwise [14]. In fact, the recent proposal [15] based on a modified Arrott-plot (MAP) [16] analysis of a magnetic system with competing interactions (hereafter, referred as nontrivial system) suggested different CEs on either side of $T_{\rm{C}}$.
Several groups have determined CEs for numerous nontrivial magnetic systems, such as ${\mathrm{La}}_{0.6}{\mathrm{Ag}}_{0.2}{\mathrm{Bi}}_{0.2}{\mathrm{MnO}}_{3}$ [7], ${\mathrm{La}}_{0.7}{\mathrm{Sr}}_{0.3}{\mathrm{MnO}}_{3}$ [17], ${\mathrm{Co}}_{3.6}{\mathrm{Fe}}_{4.4}{\mathrm{Zn}}_{8}{\mathrm{Mn}}_{4}$ [18], ${\mathrm{La}}_{0.75}{\mathrm{Sr}}_{0.25}{\mathrm{MnO}}_{3}$ [19], and ${\mathrm{Gd}}$ [20, 21, 22, 23], using the existing methodologies. It was highlighted that the reported CE values do not belong to any single UC [7, 17, 18, 19, 20]. No correlation of the reported values of CEs with the known competing interactions, eg. super-exchange, double-exchange, dipolar, Ruderman-Kittel-Kasuya-Yusida (RKKY), etc., in nontrivial systems [6, 7, 8, 9, 10, 11, 12, 13] raises a serious question: is the fundamental understanding of the critical behavior or the used methodologies or both the factors responsible for this?
It has been pointed out that the established methodologies such as iteration method (IM) [16], Kouval-Fisher plots [24], and entropy analysis [25], may yield ambiguous values of the CEs. Additionally, the determined value of $\gamma$ is essentially used to determine $\sigma$ [26] which defines the existing range of interaction in a magnetic system. The theoretical suggestions for long- and short-range interactions correspond to $\sigma\leq 1.5$ and $\sigma\geq 2.0$, respectively, often leads to conflicting interpretations, particularly when $\sigma$ falls between the two values [4, 5, 27, 28, 29, 18]. Furthermore, the $\gamma$ for a nontrivial magnetic system constitutes the effective value. So, an inaccurate determination of $\sigma$, which effectively depends on $\gamma$ [26], may result in nonphysical conclusions.
Thus, we are still left with several open problems which are necessary to resolve for better understanding of the physics of critical behavior. Some of them are as follows: There is no methodology to (i) accurately identify the range of competing interactions in nontrivial systems and their resultant effect in presence of applied magnetic field ($B=\mu_{0}H$), and (ii) systematically determine accurate and convergent values of CEs in least number of iterations. It is still unknown if the competing short- and long-range interactions constitute different UCs. Whether the two sets of exponents on either side of $T_{\rm{C}}$ are due to the intrinsic complex interactions or due to technical issues in the methodologies used for the analysis of magnetic behavior requires a thorough investigation.
@@
In this letter, we present our computational and experimental investigations on the phase transition and the critical phenomena. We report the: (i) the development of robust methodologies to determine precise and accurate values of the CEs in least number of iterations, (ii) the effect of $B$ on the nature of the competing interactions in nontrivial systems, and (iii) why the values of the CEs determined for nontrivial systems do not belong to a single UC. By performing computational investigations for trivial system, such as Ni, we first show that the iteration method may lead to ambiguous results. We then propose a method to identify the range of exchange interaction. Next, we propose new methodologies using the computationally simulated magnetization isotherms to determine correct and convergent values of the CEs in least number of iterations. Subsequently, we show that magnetic systems hold identical values of the CEs on either side of $T_{\rm{C}}$. Finally, by performing experimental investigations for single crystals of Gd, we show that the CEs for nontrivial systems do not belong to a single UC due to the competing effect of the existing exchange interactions.
@@
Figure 1: (a) The temperature-dependent magnetization ($M-T$) plot obtained from the Vampire atomistic simulation for monatomic Ni. Inset of (a) represents the first derivative of the $M-T$ resulting $T_{\rm{C}}=655$ K. (b) Field-dependent magnetization ($M-H$) estimated at $T=T_{\rm{C}}$. Inset of (b) represents the linear fit to log-log plot of $M-H$. (c) The modified Arrott-plots constructed using the critical exponents determined from standard procedure. (d) Normalized slope ($NS$) vs. $T$ plots constructed using the CEs of the standard UCs, and the CEs determined from IM, MIM, SP and new proposal (NP). (e) The plots of $M_{S}$ and $\chi_{0}^{-1}$ against $T$. Symbols represent simulated data while solid lines are fit to respective curves. (f) intercept ($I$) vs. $T$ plots constructed using the CEs determined using SP and of standard UCs. (g) $I$ vs. $T$ plots for different values of $\beta$ and $\gamma$ which are related as $\gamma=\beta\times(\delta-1)$. (h) Variation of $NS_{\rm{avg}}$ with $n$ which is related with $\beta$ as $\beta(n)=\beta_{i}+\Delta\beta\times n$. Inset of (h) represent the $NS_{\rm{avg}}$ vs. $\beta$ plots constructed using few sets of $\beta$ and corresponding $\gamma$ as mentioned in (g). MF: mean field, TMF: tricritical MF and H: Heisenberg.
@@
In order to determine precise values of CEs, we first proceed with the analysis of the magnetization ($M$) data for Ni. The $T$-dependent magnetization ($M-T$) for both zero- and applied-$B$ are determined by carrying out atomistic simulations based on solution of Landau-Lifshitz-Gilbert equation using Vampire simulation package [30, 31]. $T_{\rm{C}}=655$ K is determined from the minima of the first-derivative [inset of Fig. 1(a)] of the $M$ vs. $T$ plot [Fig. 1(a)]. Using the relations, $M\propto|\epsilon|^{\beta}$, $M\propto(\mu_{0}H)^{\frac{1}{\delta}}$ at $T=T_{\rm{C}}$ and the Widom relation $\delta=1+\frac{\gamma}{\beta}$  [5, 4, 32, 33], we find $\beta=0.344(2)$ [Fig. 1(a)], $\delta=4.822(2)$ [Fig. 1(b)] and $\gamma=1.315(8)$, respectively. Here, $\epsilon=\frac{T-T_{\rm{C}}}{T_{\rm{C}}}$ is the reduced $T$. We denote the above methods of determination of CEs as the standard procedure (SP). We further define the normalized slopes ($NSs$) as the ratio of slope of the MAP at a given $T$ and that at $T_{\rm{C}}$, see [Fig. 1(c)]. $NSs\sim 1$ thus obtained for $T\leq T_{\rm{C}}$ [Fig. 1(d)] [31] confirms the reliability of the obtained values of CEs that belongs to 3D XY UC [34, 35, 36, 37, 38]. However, the values of CEs obtained by IM, viz., $\beta=0.362(1)$ and $\gamma=1.222(10)$, are different [Fig. 1(e)] suggesting that IM  [31] is not yielding the coorect values of the CEs. Moreover, the values of $NSs$, plotted using CE values obtained by IM, show deviation from 1 on either side of $T_{\rm{C}}$, demonstrating further that IM is an incorrect approach to determine CEs. Additionally, the CEs determined using MIM for $T\leq T_{\rm{C}}$ shows 3D XY type short-range ($\sigma\approx 2$) [31] interaction and $NS\sim 1$ [18]. However, appropriate values of CEs for $T\geq T_{\rm{C}}$ could not be determined using MIM since $NSs$ do not come close to 1 for the used $T$-range. Further, we note that in many real systems, SP may not be always applicable for critical analysis, see discussions below. Therefore, we propose novel methodologies for critical analysis as follows.
@@
Figure 1(f) shows the plots of the $T$-dependence of intercepts obtained from the linear fits for MAPs corresponding to different UCs. Remarkably, we find that the intercepts estimated by both SP as well as by using the CEs of short-range standard UCs show a crossover at $T\sim T_{\rm{C}}$. Interestingly, if CEs are estimated considering MF UC representing a long-range interaction, the crossover $T$ ($T_{2}=669$ K) is found to deviate from $T_{\rm{C}}$. Thus, from the detailed analysis of CEs for Ni, we identify the approach of intercept analysis as a reliable method to determine the short-range interaction in Ni for which we computationally determine the $M-H$ isotherms by considering anisotropy value as zero. The same approach may also be applicable to identify long-range interactions if the intercepts, determined using the CEs of MF UC, show a crossover around $T_{\rm{C}}$ (see discussions below). This suggests that the new method of intercept analysis may be a proper approach to determine the range of interactions. By using this approach, we demonstrate (see below) the role of itinerant electrons in the magnetic behavior of Gd. Intriguingly, we note further that the different values of $\beta$ and $\gamma$ corresponding to short-range standard UCs lead to the same value of $\delta$ which is equal to the $\delta$ [Fig. 1(b)] as determined from the $M-H$ at $T_{\rm{C}}$. Additionally, the values of the intercepts determined using the CEs of 3D Ising, 3D XY, 3D Heisenberg and SP converge to 00 at $T=T_{\rm{C}}$ (the crossover $T$). This suggests that the values of CEs related with $\delta$ may lead to value of intercept $\approx 0$ at $T=T_{\rm{C}}$ for multiple set of $\beta$ and $\gamma$.
@@
Using the above information, we now describe our new proposal for critical analysis as follows. We vary $\beta$ in a controlled way as $\beta(n)=\beta_{i}\pm\Delta\beta\times n$, where $\beta_{i}$ is the initially selected value of $\beta$, $\Delta\beta$ is the increment in $\beta$ for each iteration and $n$ is the iteration number. $\beta_{i}$ can be chosen as belonging to a standard UC or a random value with the condition that the chosen set of $\beta$ and corresponding $\gamma$ lead to quasi-linear behavior of MAPs. The corresponding values of $\gamma$ are determined for each $\beta(n)$ using the Widom relation [33] as $\gamma=\beta(n)\times(\delta-1)$, where $\delta$ is directly obtained from $M-H$ curve at $T=T_{\rm{C}}$ [Fig.  1(b)]. The intercepts of MAPs thus obtained are found to exhibit a crossover at $T_{\rm{C}}$ [Fig. 1(g)] for multiple sets of $\beta$ and $\gamma$ values. In order to determine appropriate values of $\beta$ and $\gamma$, the MAPs obtained as above are used to estimate an average value ($NS_{\rm{avg}}$) of $NSs$ for $T\leq T_{\rm{C}}$ and $T\geq T_{\rm{C}}$ separately. $NS_{\rm{avg}}$ is defined by $NS_{\rm{avg}}=\frac{\sum_{i=1}^{N}(NS)_{i}}{N}$, where $i$ denotes the $M-H$ isotherm at a given $T$ and $N$ is the total number of the isotherms appearing in the calculation. The $NS_{\rm{avg}}$ thus determined for each iteration $n$ are plotted in Fig. 1(h). According to our proposal, the set of values of $\beta$, $\delta$ and $\gamma$ corresponding to $NS_{\rm{avg}}=1$ are the appropriate values of the CEs for a system. Remarkably, using our proposal for $T\leq T_{\rm{C}}$, we determine $\beta=0.343(3)$ and $\gamma=1.311(18)$ which are consistent with the CEs obtained using the SP. The error in $\beta$ is manually controllable while the error in $\gamma$ depends on the imposed error in $\beta$ and the error appearing in $\delta$. In the above proposal, we vary $\beta$ and $\gamma$ in iterative way. However, we propose another unconventional and much simpler approach to determine appropriate values of CEs as follows [31]. Select at least three sets of $\beta$ and corresponding $\gamma$ such that the estimated $NS_{\rm{avg}}$ falls within $1\pm 0.05$. One can extend or reduce the above defined range for $NS_{\rm{avg}}$ within the range of linear variation of the $NS_{\rm{avg}}$ vs. $\beta$ plot [31], see the inset of Fig. 1(h). Thus, using the $intercept$ and $slope$, determined from the linear fit to $NS_{\rm{avg}}$ vs. $\beta$ plot, one can directly estimate $\beta=\frac{1-intercept}{slope}$ and $\gamma=\frac{1-intercept}{slope}\times(\delta-1)$ [31]. The $\beta=0.344(3)$ and $\gamma=1.315(18)$) thus obtained are the same as the values of CEs determined using SP.
@@
However, for $T\geq T_{\rm{C}}$, we find that the values of CEs ($\beta=0.402(5)$ and $\gamma=1.533(20)$) resulting $NS_{\rm{avg}}\sim 1$ are significantly different than as obtained for $T\leq T_{\rm{C}}$, see Fig. 1(h). As our new approach shows that for $T\leq T_{\rm{C}}$ the CE values are consistent with that obtained by the SP and NS $\sim$ 1, whereas for $T\geq T_{\rm{C}}$, the NS values are significantly different from 1 [Fig. 1(d)], it suggests that the CEs obtained for $T\geq T_{\rm{C}}$ do not define the actual critical behavior. This is further supported from $\beta$ which is undefined for above $T_{\rm{C}}$ [32]. We infer that the consideration of identical range of $T$ for below and above $T_{\rm{C}}$ may have resulted to such discrepancy. In the following, we verify this and discuss our proposal for the estimation of appropriate range of $T$ and $B$ for below and above $T_{\rm{C}}$ for critical analysis.
@@
Figure 2: (a) The field-dependent variation of $\delta$ estimated for different applied magnetic field ($B$) ranges as mentioned. (b) and (c) represent the variation of $\beta$ and $\gamma$, respectively, for different $B$ and temperature ($T$) ranges determined using our new proposal (NP). The $T$ ranges are mentioned in below Figs. (d)--(i). (d)--(f) and (g)--(i) represent NS and intercept (I) against $T$ plots, respectively, constructed using the CEs obtained using NP for different $B$ and $T$ ranges.

### v2

# New Methods for Critical Analysis: Revealing the Simultaneous Existence of Universality Classes in Nontrivial Magnetic Systems

@@ Abstract
In magnetic systems, the microscopic constituents exhibit power law behavior near the paramagnetic transition temperature, $T_{\rm{C}}$. The critical exponents (CEs) associated with the physical quantities that demonstrate singular behavior at $T_{\rm{C}}$ illustrate the critical behavior, specifically the range and type of exchange interactions emerging in magnetic systems. However, it is realized that the developed methodologies may not yield accurate values of CEs, especially for magnetic systems with competing interactions, referred to as nontrivial magnetic systems. Currently, no comprehensive method effectively addresses the competing effects of the range of magnetic interactions among the constituent entities emerging in such systems. Additionally, there is no definitive explanation for CE values that do not belong to any single universality class. Here, we present new methodologies for critical analysis aimed at determining both the range of exchange interaction(s) and appropriate values of CEs. Using computational and experimental investigations, we analyze the magnetic behavior of trivial Ni and nontrivial Gd. Our findings demonstrate that (i) the critical behavior remains the same on either side of $T_{\rm{C}}$, (ii) the critical behavior associated with local electron moments remains unaffected by the magnetic field, and (iii) in Gd, the critical role of competing interactions becomes evident: local electron moments follow a three-dimensional Ising-type short-range interaction, while itinerant electron moments exhibit a mean-field-type long-range Ruderman–Kittel–Kasuya–Yosida (RKKY) interaction, which weakens under an external magnetic field due to the localization effect on itinerant electrons.
@@
In the studies of phase transitions for cases, such as condensation of gases, melting of solids, magnetic ordering in solids, etc., the understanding of the behavior of the systems adjacent to the critical point is of fundamental interest [1, 2, 3]. Generally, the physics of phase transition is governed by the behavior of the microscopic constituents which interact strongly in a cooperative fashion. Critical exponents (CEs), which define the power law behavior of physical quantities exhibiting singularities at the critical points, indicate the qualitative nature, i.e., the critical behavior of a system [4, 5]. The types of cooperative interactions (short-range, long-range, etc.) among the constituent entities play a significant role in defining the CEs. In magnetic systems, the magnetic moments are known to interact typically via different types of exchange interactions resulting in a set of CEs [6, 7, 8, 9, 10, 11, 12]. Different methodologies have been suggested for the determination of a set of CEs ($\alpha,\beta,\gamma$, $\delta$, etc.) corresponding to physical quantities such as specific heat ($C_{\rm{v}}$), spontaneous magnetization ($M$), susceptibility ($\chi$) and magnetization isotherm recorded at the paramagnetic transition temperature ($T_{\rm{C}}$), respectively for various universality classes (UCs), viz., Heisenberg, XY, Ising, mean field (MF), or tricritical MF [4, 5]. The CEs are mutually related and they depend critically on the dimensionality (of space and spin) and the range of microscopic interactions.
Several experimental methodologies have been developed to investigate the critical behavior of magnetic systems [13, 14, 15]. Although, most methodologies suggest same CEs for two sides of a transition, a recent theoretical work suggested otherwise [16]. In fact, the recent proposal [17] based on a modified Arrott-plot (MAP) [13] analysis, (i) pointed out that the existing methodologies for critical analysis may yield ambiguous values of CEs, (ii) proposed a new method for critical analysis, and (iii) supported the claim of ref. [16]. Although many important points have been settled for the study of critical behavior, we are still left with several crucial open problems, as discussed below, which are needed to be resolved to fully understand the critical behavior of magnetic systems.
The critical behavior of numerous nontrivial magnetic systems, such as ${\mathrm{La}}_{0.6}{\mathrm{Ag}}_{0.2}{\mathrm{Bi}}_{0.2}{\mathrm{MnO}}_{3}$ [7], ${\mathrm{La}}_{0.7}{\mathrm{Sr}}_{0.3}{\mathrm{MnO}}_{3}$ [18], ${\mathrm{Co}}_{3.6}{\mathrm{Fe}}_{4.4}{\mathrm{Zn}}_{8}{\mathrm{Mn}}_{4}$ [19], ${\mathrm{La}}_{0.75}{\mathrm{Sr}}_{0.25}{\mathrm{MnO}}_{3}$ [20], and ${\mathrm{Gd}}$, etc. [21, 22, 23, 24] have been reported in literature. Here, the magnetic systems with competing interactions are referred to as nontrivial. In most of the cases, it was highlighted that the reported CE values in nontrivial magnetic systems do not belong to any single UC [7, 18, 19, 20, 21] which is a serious impediment in the understanding of critical behavior of magnetic systems in general. The answer to this observation remains elusive; therefore, a deeper understanding of critical behavior in nontrivial magnetic systems is crucial and requires more thorough investigation. Moreover, the value of $\gamma$ obtained using the developed methodologies is essentially used to determine $\sigma$ [25] which defines the range of interaction in magnetic systems. The theoretical suggestions for long- and short-range interactions correspond to $\sigma\leq 1.5$ and $\sigma\geq 2.0$, respectively, which often leads to conflicting interpretations, particularly when $\sigma$ falls between the two values [4, 5, 26, 27, 28, 19]. We note here that the $\gamma$ estimated for nontrivial magnetic systems is an effective value. Therefore, an inaccurate determination of $\sigma$- which effectively depends on $\gamma$ [25] -- is likely to lead to nonphysical conclusions. This highlights the lack of a reliable methodology for accurately identifying the range of competing interactions, especially in nontrivial systems. The question of whether the two sets of exponents on either side of $T_{\rm{C}}$ are due to the intrinsic complex interactions or due to technical issues in the methodologies used for the analysis of magnetic behavior requires a thorough investigation. Furthermore, the influence of external magnetic field ($B=\mu_{0}H$) on the competing interactions remains an open problem to investigate. On the technical front, a new methodology is a prerequisite to determine the accurate values of CEs in least number of iterations.
@@
In this letter, we present our detailed computational and experimental investigations of the critical behavior that potentially emerge in trivial as well as nontrivial magnetic systems. For computational investigations, we chose Ni as a trivial system. (i) We report a novel methodology to determine precise and accurate values of the CEs in least number of iterations. Using the proposed method, robust and convergent results are obtained for Ni. (ii) We propose a new methodology to qualitatively identify the range of exchange interaction in these systems. After confirming the correctness of the results for Ni using our developed methodologies, we investigate the critical behavior of nontrivial single crystalline Gd. (iii) With our new approach of analysis of experimentally determined magnetization isotherms, we identify that the local and itinerant electron moments in Gd exhibit different types of exchange interactions following different UCs thereby resulting in effective values of CEs. (iv) We also analyze the effect of $B$ on the nature of the competing interactions in nontrivial systems. (v) Subsequently, we show that magnetic systems hold identical values of the CEs on either side of $T_{\rm{C}}$.
@@
Figure 1: (a) The $M-T$ plot obtained using the Vampire atomistic simulation for Ni. Inset of (a) represents the first derivative of the $M-T$ resulting $T_{\rm{C}}=655$ K. (b) The $M-H$ isotherm estimated at $T=T_{\rm{C}}$. Inset shows the linear fit to log-log plot of $M-H$. (c) The MAPs constructed with CEs determined using SP. (d) $NS$ vs. $T$ plots constructed using the CEs of the standard UCs, and the CEs determined from IM, MIM, SP and the new proposal (NP). (e) The plots of $M_{S}$ and $\chi_{0}^{-1}$ against $T$. Symbols represent simulated data while solid lines are fit to respective curves. (f) $I$s vs. $T$ plots constructed using the CEs determined using SP and of standard UCs. (g) $I$s vs. $T$ plots for different values of $\beta$ and $\gamma$ which are related as $\gamma=\beta\times(\delta-1)$. (h) Variation of $NS_{\rm{avg}}$ with $n$ which is related with $\beta$ as $\beta(n)=\beta_{i}+\Delta\beta\times n$. Inset of (h) represent the $NS_{\rm{avg}}$ vs. $\beta$ plots constructed using a few sets of $\beta$ and corresponding $\gamma$ as mentioned in (g). MF: mean field, TMF: tricritical MF and H: Heisenberg.
@@
In order to determine precise values of CEs, we first proceed with the analysis of the magnetization ($M$) data for Ni. The temperature ($T$)-dependent magnetization ($M-T$) for both zero- and applied-$B$ are determined by carrying out atomistic simulations based on solution of Landau-Lifshitz-Gilbert equation using Vampire simulation package [29, 30]. $T_{\rm{C}}=655$ K is determined from the minima of the first-derivative [inset of Fig. 1(a)]. Using the relations, $M\propto|\epsilon|^{\beta}$, $M\propto(\mu_{0}H)^{\frac{1}{\delta}}$ at $T=T_{\rm{C}}$ and the Widom relation $\delta=1+\frac{\gamma}{\beta}$  [5, 4, 34, 35], we obtain $\beta=0.344(2)$ [Fig. 1(a)], $\delta=4.822(2)$ [Fig. 1(b)] and $\gamma=1.315(8)$, respectively, which corresponds to 3D-XY UC. Here, $\epsilon=\frac{T-T_{\rm{C}}}{T_{\rm{C}}}$ is the reduced $T$. The above methods of determination of CEs are identified as the standard procedure (SP). We note here that in many real systems, SP may not be always applicable for critical analysis, see discussions below. In the following, we analyze the primary issues involved with the other existing methods for critical analysis such as IM, modified IM (MIM) [17], etc. to determine the critical behavior of magnetic systems.
A. Arrot and J. E. Noakes [13] suggested a linear relationship between $M^{1/{\beta}}$ and $(B/M)^{1/{\gamma}}$ for appropriate values of CEs. The plots of $M^{1/{\beta}}$ vs. $(B/M)^{1/{\gamma}}$ is termed as MAPs, see Fig. 1(c). The linearity of the MAPs plotted using the values of $\beta$ and $\gamma$ can be checked by determining the normalized slopes ($NS$s), defined as slopes of the MAP at a given $T$ normalized with respect to that at $T=T_{\rm{C}}$. Figure 1(d) shows that the values of $NS$, if determined using the correct CE values obtained by the SP, are $\sim 1$ for $T\leq T_{\rm{C}}$. We, therefore, use $NS\sim 1$ for $T\leq T_{\rm{C}}$ as the criterion to test the validity of a method of critical analysis. We note here that for $T>T_{\rm{C}}$, the values of $NS$ deviate significantly from $1$ which we relate to choosing the appropriate range of $T$ for the analysis of critical behavior of a system  [17, 30], see analysis and discussion on this below.
In order to check the validity of IM for critical analysis, we determine the CE values for Ni using the IM. The values of CEs, viz. $\beta=0.362(1)$ and $\gamma=1.222(10)$ [Fig. 1(e)] obtained by IM are markedly different from the ones obtained using SP. Moreover, as shown in Fig. 1(d), the $NS$s for MAPs plotted using these values of $\beta$ and $\gamma$ deviate significantly from the criterion $NS\sim 1$ on either side of $T_{\rm{C}}$ demonstrating that the IM may not be an appropriate approach for critical analysis of a magnetic system. Interestingly, the CEs ($\beta=0.343(3)$, $\gamma=1.309(8)$) determined using MIM and $NS\sim 1$ for $T\leq T_{\rm{C}}$ shows 3D XY type short-range ($\sigma\approx 2$) interaction [19, 30]. However, appropriate values of CEs for $T\geq T_{\rm{C}}$ could not be determined using MIM as $NS$ values deviate significantly from $1$, see [30]. Thus, it is clear that the existing methodologies may not provide robust results which has a profound effect on the general understanding of phase transitions of a system. Below, we propose novel methodologies for the analysis of critical behavior of magnetic systems and demonstrate with examples of Ni and Gd that the criterion we discussed above is fully satisfied.
@@
Figure 1(f) shows the plots of the $T$-dependence of intercepts ($I$s) obtained from the linear fits of MAPs for Ni corresponding to different UCs. It is clear that the $I$s, estimated by both SP as well as by using the CEs of short-range standard UCs, show a crossover at $T\sim T_{\rm{C}}$. However, if $I$s are estimated considering MF UC representing a long-range interaction, the crossover $T$ ($T_{2}=669$ K) deviates from $T_{\rm{C}}$. Since the short-range magnetic interaction is established for Ni from the SP of critical analysis, we identify the crossover of the $I$s at $T=T_{\rm{C}}$ as the criterion to determine the range of interaction in a magnetic system. The same approach may also be applicable to identify long-range interactions if the $I$s, determined using the CEs of MF UC, show a crossover around $T_{\rm{C}}$. We show below that this new method of analysis based on $I$s of MAPs may be a proper approach to determine the appropriate range of interactions. By using this approach, we have demonstrated below the role of itinerant electrons in the magnetic behavior of Gd. Moreover, we note that the different sets of $\beta$ and $\gamma$ values corresponding to different short-range standard UCs lead to the same value of $\delta$ as determined from the $M-H$ at $T_{\rm{C}}$, see Fig. 1(b). Additionally, the values of the $I$s of MAPs determined using the CEs corresponding to 3D Ising, 3D XY, 3D Heisenberg and SP converge to 00 at $T=T_{\rm{C}}$ (the crossover $T$). Thus, although the type of interaction (i.e., short-range) has been identified, the exact nature of short-range interaction is still to be determined.
@@
To determine the accurate values of the CEs, we proceed as follows. We vary $\beta$ in a controlled way such that $\beta(n)=\beta_{i}\pm\Delta\beta\times n$, where $\beta_{i}$ is the initially selected value, $\Delta\beta$ is the increment in $\beta$ for each iteration and $n$ is the iteration number. To start with, $\beta_{i}$ can be chosen as a theoretical value corresponding to any standard UC or even any random value with the condition that the chosen set of $\beta$ and corresponding $\gamma$ lead to linear behavior of MAPs, see Fig. 1(c) for example. The corresponding values of $\gamma$ are determined for each $\beta(n)$ using the Widom relation [35], $\gamma=\beta(n)\times(\delta-1)$, where $\delta$ is directly obtained from $M-H$ curve at $T=T_{\rm{C}}$ [Fig.  1(b)]. The $I$s of MAPs thus obtained are found to exhibit a crossover at $T_{\rm{C}}$ [Fig. 1(g)] for multiple sets of $\beta$ and $\gamma$ values. In order to determine appropriate values of $\beta$ and $\gamma$, the MAPs obtained as above are used to estimate an average value ($NS_{\rm{avg}}$) of $NSs$ for $T\leq T_{\rm{C}}$ and $T\geq T_{\rm{C}}$ separately. $NS_{\rm{avg}}$ is defined by $NS_{\rm{avg}}=\frac{\sum_{i=1}^{N}(NS)_{i}}{N}$, where $i$ denotes the $M-H$ isotherm at a given $T$ and $N$ is the total number of the isotherms used in the analysis. The $NS_{\rm{avg}}$ thus determined for each iteration $n$ are plotted in Fig. 1(h). According to our proposal, the set of CE values ($\beta$, $\delta$ and $\gamma$) corresponding to $NS_{\rm{avg}}=1$ and $NS\sim 1$ are the appropriate values of the CEs for a system. Using our method proposed here, we determine $\beta=0.343(3)$ and $\gamma=1.311(18)$ for $T\leq T_{\rm{C}}$ which are remarkably consistent with the CEs obtained using the SP. The error in $\beta$ is manually controllable while the error in $\gamma$ depends on the imposed error in $\beta$ and the error appearing in $\delta$.
@@
However, for $T\geq T_{\rm{C}}$, we find that the values of CEs ($\beta=0.402(5)$ and $\gamma=1.533(20)$), resulting $NS_{\rm{avg}}\sim 1$, are significantly different than as obtained for $T\leq T_{\rm{C}}$, see Fig. 1(h). Moreover, although $NS_{\rm{avg}}\sim 1$ for $\beta=0.402(5)$, the NS values show significant deviation from the criterion of $NS\sim 1$ [Fig. 1(d)]. This suggests that the CEs obtained for $T\geq T_{\rm{C}}$ may not define the critical behavior. This is further supported from $\beta$ which is undefined for above $T_{\rm{C}}$ [34]. Moreover, it is difficult to identify the $B$-dependent $T$-range up to which the short-range exchange interaction survives since the thermal fluctuation mainly control the physics above $T_{\rm{C}}$ in the vicinity of $T_{\rm{C}}$. The later statement indicate that one may get different values of $\gamma$ for below and above $T_{\rm{C}}$ depending on the properties of the chosen system for investigation as suggested in ref. [16]. This is because the correlation length $\xi$ is usually different for below and above $T_{\rm{C}}$ in nontrivial magnetic systems. From our detailed analysis, we infer that the consideration of identical range of $T$ for the analysis below and above $T_{\rm{C}}$ may have resulted in such a discrepancy. In the following, we verify this and discuss a procedure to estimate the appropriate range of $T$ and $B$ to be used for critical analysis for below and above $T_{\rm{C}}$.
@@
Figure 2: (a) The field range dependence of $\delta$. Different field ranges are mentioned in the Figure. (b) and (c) represent the variation of $\beta$ and $\gamma$, respectively, for different $B$ and $T$ ranges determined using our new proposal (NP). (d)--(f) and (g)--(i) represent $NS$ and $I$ vs. $T$ plots, respectively, constructed using the CEs obtained using NP for different $B$ and $T$ ranges.